Find general solution of equation 2x(y+z^2)p+y(2y+z^2)q=z^3 and deduce that yz(z^2+yz-2y)=x^2 is a solution
1
Expert's answer
2020-08-10T17:54:26-0400

"Solution"

"2x(y+z^2)\frac{dz}{dx}+y(2y+z^2)\frac{dz}{dy}=z^2"


"\frac{dx}{2x(y+z^2)}=\frac{dy}{y(2y+z^2)}=\frac{dz}{z^2}"



"y\frac{dx}{2xy^2+2xyz^2}=x\frac{dy}{2xy^2+xyz^2}=xy\frac{dz}{xyz^2}"


"\frac{ydx−xdy−xydz}{2xy2+2xyz2−2xy2−xyz2−xyz2}=\frac{ydx−xdy−xydz}{0}"

"ydx\u2212xdy\u2212xydz=0"


"\frac{dx}{x}−\frac{dy}{y}−dz=0"

"lnx\u2212lny\u2212z=c"

"z=ln( \frac{x}{y})+C"

2.)


"yz(z^2+2z\u22122y)=x^2"


"yz^3+2yz^2\u22122y2z=x^2"

Differentiate respect to x:


"3yz^2p+4yzp\u22122y^2p=2x"

"py(3z^2+4z\u22122y)=2x"

Differentiate respect to y:


"z^3+3z^2yq+2z^2+4yzq\u22124yz\u22122y^2q=0"

"qy(3z^2+4z\u22122y)+z^3+2z^2\u22124yz=0"

"\therefore" we have:


"3z^2+4z−2y=\frac{2x}{py}"


"\frac{2xq}{p}+z^3+2z^2−4yz=0"

Since


"p=\frac{1}{x}"

"q=\frac{−1}{y}"

then:


"\frac{−2x^2}{y}+z^3+2z^2−4yz=0"

"\u22122x^2+y(z^3+2z^2\u22124yz)=0"

"yz(z^2+2z\u22124y)=2x^2"

So:


"\frac{z^2+2z−4y}{z^2+2z−2y}=2"

"z^2+2z\u22124y=2z^2+4z\u22124y"

"z^2+2z=0"

So we get that the statement can be proved if

"z=0, or"

"z=\u22122"